# MMHMR: Generative Masked Modeling for Hand Mesh Recovery

### v1

@@ 4 Experiments
MMHMR Ours 61.0 87.1 94.8 62.1 90.2 95.0 59.3 88.3 94.4
Occluded Joints
FrankMocap [50] ICCVW 2021 9.2 28.0 46.9 11.0 33.0 55.0 8.4 26.9 45.1
METRO [38] CVPR 2021 7.0 23.6 42.4 10.2 32.4 53.9 8.0 26.2 44.7
MeshGraphormer [39] ICCV 2021 7.9 25.7 44.3 10.9 33.3 54.1 9.3 32.6 51.7
HaMeR [46] CVPR 2024 27.2 60.8 78.9 25.9 60.8 80.7 23.0 56.9 76.3
MMHMR Ours 29.4 64.1 80.3 31.4 64.2 83.6 29.4 65.3 80.1
Datasets.To train the hand pose tokenizer, we employed a diverse set of datasets to capture a wide range of hand poses and interactions. This includes DexYCB [9], InterHand2.6M [43], MTC [61], and RHD [73]. For training MMHMR, we utilized a diverse dataset, following a similar setup as in [46] to ensure a fair comparison. Specifically, the training data was drawn from FreiHAND [74], HODv2 [25], MTC [61], RHD [73], InterHand2.6M [43], H2O3D [25], DexYCB [9], COCO-Wholebody [32], Halpe [20], and MPII NZSL [51].
Evaluation Metrics. Following standard protocols [46, 37, 71], MMHMR evaluated on reconstructed 3D joints using PA-MPJPE and AUCJ, while 3D mesh vertices were evaluated with PA-MPVPE, AUCV, F@5mm, and F@15mm. Additionally, to examine MMHMR’s generalization and accuracy in diverse real-world settings, we employed the Percentage of Correct Keypoints (PCK) [46] metric at multiple thresholds, ensuring a robust evaluation of its performance across varied conditions. To evaluate hand image generation with mesh-guided control, we compute FID-H and KID-H on cropped hand regions, following previous works [44] and use MediaPipe [65] as a hand detector to measure confidence.
3D Reconstruction Accuracy Evaluation.To comprehensively evaluate MMHMR’s 3D joints and mesh reconstruction capabilities, we used the HO3Dv3 [26] and FreiHAND datasets. HO3Dv3 [26], the largest 3D hand pose test dataset with 20K annotated images, serves as a rigorous test ground with diverse hand poses and complex scenarios. MMHMR was directly tested on HO3Dv3 without prior training on this dataset, assessing its ability to generalize to new data and confirming the model’s robustness in challenging settings. This evaluation underscores MMHMR’s resilience to dataset-specific biases, proving its adaptability across varied scenarios. Additionally, we evaluated MMHMR on the FreiHAND dataset [74], which includes 4K images spanning controlled environments. For fair comparison, as previous methods like MeshGraphormer [39] and HHMR [37] were evaluated on FreiHAND, we evaluated MMHMR trained solely on the FreiHAND [74].
Real-World Robustness Evaluation. To validate MMHMR’s adaptability, we tested it on the HInt benchmark [46] without prior training. HInt introduces diverse real-world conditions, including varied lighting, angles, and hand-object interactions, offering a realistic evaluation framework. Using the Percentage of Correct Keypoints (PCK) metric, we assessed MMHMR on HInt-NewDays [12], HInt-EpicKitchensVISOR [15], and HInt-Ego4D [24], highlighting its robust generalization and effectiveness in complex environments.
4.1 Comparison to State-of-the-art Approaches
We evaluate MMHMR against a range of state-of-the-art methods (SOTA) on the HO3Dv3 [26], FreiHAND [74], and HInt [46] benchmarks, as detailed in Table 1, Table 2, and Table 3. MMHMR consistently outperforms competing methods across key evaluation metrics, demonstrating robust accuracy in 3D hand reconstruction. Notably, we perform zero-shot evaluations on both the HO3Dv3 and HInt benchmarks to assess MMHMR’s generalizability. A core contributor to MMHMR’s success is its capability to model and refine uncertainty, making it highly effective in scenarios with complex hand poses and significant occlusions. On the HO3Dv3 dataset (Table 1), MMHMR achieves a PA-MPJPE reduction of approximately 19.5% and a PA-MPVPE reduction of 15.7% compared to the existing SOTA method. This improvement underscores MMHMR’s precision in handling challenging hand poses and occlusions. Similarly, on the HInt benchmark (Table 3), MMHMR achieves notable improvements in occluded joint reconstruction at the PCK@0.05 threshold on the HInt benchmark. Specifically, MMHMR shows an 8.1% increase on HInt-NewDays, a 21.2% increase on HInt-VISOR, and a 27.8% increase on HInt-Ego4D compared to the closest SOTA methods. These gains underscore MMHMR’s effectiveness in synthesizing unobserved parts and handling real-world occlusions with high accuracy. This highlights MMHMR’s strength in synthesizing unobserved parts, enabling accurate reconstruction of occluded regions.
Mesh-Guided Control for Hand Image Generation. As shown in the Generation Pipeline (Figure 4) and Table 13, MMHMR demonstrates adaptability in generating realistic hand images under diverse conditions. We train ControlNet [66]on DexYCB [9] and test it on HO3Dv3 [26]. MMHMR processes input images through a mesh estimator i.e. (MMHMR) to extract 3D hand meshes, which serve as control signals in ControlNet [66] alongside text prompts. Compared to HaMer [46] as a mesh estimator method [46], MMHMR achieves lower FID-H and KID-H scores, indicating better image quality, and a higher hand detection confidence score, reflecting improved anatomical accuracy. Accurate mesh estimation ensures precise control signals for consistent, high-quality hand generation. More details are in the supplementary material.
Table 4: Mesh-Guided Control for Hand Image Generation
Mesh Estimator FID-H $\downarrow$ KID-H $\downarrow$ Hand Det Conf. $\uparrow$
HaMer [46] 45.65 0.035 0.865
MMHMR (Ours) 40.23 0.027 0.912
Figure 4: Generation Pipeline: MMHMR extracts 3D hand meshes from images during training, using them with text prompts to guide ControlNet. At inference, only text and mesh signals are needed to generate realistic hand images.
4.2 Ablation Study
The effectiveness of MMHMR is grounded in its mask modeling and iterative decoding techniques. This ablation study examines how iterative refinement and mask-scheduling strategies impact model performance. Additional experimental results and visualizations in the Supplementary Material provide in-depth analyses of the key factors driving MMHMR’s performance. These include (1) architectural components, such as the VQ-MANO tokenizer, GGPM design, Context-Infused Masked Synthesizer, and feature resolution choices, (2) training strategies, including mask scheduling and regularization via keypoint and MANO losses, and (3) the model’s limitations.
Effectiveness of Proposed Components. Our ablation study on the HO3Dv3 dataset [26] (Table 5) underscores the importance of MMHMR’s components. The Upsampler plays a key role in enhancing fine-grained details and resolution, while the 2D OpenPose Context provides essential spatial cues; their absence leads to reduced accuracy. The x-Attention Head is crucial for maintaining hand shape stability by integrating shape and camera parameters, with its removal resulting in a decline in 3D accuracy. Iterative Decoding is vital for refining predictions by leveraging contextual cues, with its absence causing significant performance degradation. The largest performance drop is observed when Graph-Guided Pose Modeling (GGPM) is removed, emphasizing its critical role in capturing joint dependencies and ensuring anatomical coherence.
Method PA-MPJPE (↓) PA-MPVPE (↓) F@5mm (↑) AUCJ (↑) AUCV (↑)
@@ 4.2 Ablation Study
w/o. x-Attention Head 7.3 7.1 0.655 0.855 0.858
w/o. GGPM 7.3 7.3 0.645 0.853 0.854
w/o. Iterative Decoding 7.2 7.2 0.654 0.857 0.857
MMHMR (Full) 7.0 7.0 0.663 0.860 0.860
Impact of Iterative Confidence-Guided Sampling. The number of iterations in confidence-guided sampling impacts both accuracy and efficiency in 3D hand pose estimation. As shown in Table 6, increasing iterations from 1 to 5 improves accuracy, with PA-MPVPE dropping from 7.2 to 7.0 on HO3Dv3 and from 5.8 to 5.5 on FreiHAND. This refinement enhances precision by leveraging contextual cues and 2D pose guidance. However, increasing to 10 iterations shows minimal benefit, making 5 iterations optimal for balancing accuracy and computational cost.
Table 6: Iterations in Iterative Confidence-Guided Sampling
HO3Dv3 FreiHAND
# of iter. PA-MPJPE PA-MPVPE PA-MPJPE PA-MPVPE
1 7.2 7.2 5.9 5.8
3 7.1 7.1 5.8 5.7
5 7.0 7.0 5.7 5.5
10 7.1 7.0 5.7 5.5
Masking Ratio during Training. The ablation study in Table 7 shows that a broader masking range $\gamma(\tau\in\mathcal{U}(0,0.7))$ achieves optimal results on HO3Dv3 and FreiHAND, with the lowest PA-MPVPE values of 7.0 and 5.5, respectively. This cosine-based masking strategy, where the model learns to reconstruct from partially masked sequences, enhances robustness in 3D hand reconstruction. Narrower masking ranges, such as $\gamma(\tau\in\mathcal{U}(0,0.3))$, increase error, highlighting the importance of challenging the model with broader masking for better generalization.
@@
5 Conclusion
In this paper, we introduced MMHMR, a novel generative masked model designed for accurate and robust 3D hand mesh reconstruction from single RGB images. MMHMR addresses the longstanding challenges posed by complex hand articulations, self-occlusions, and depth ambiguities by leveraging a generative framework that captures and refines hand pose distributions. Central to our approach are two key components: VQ-MANO, which encodes 3D hand articulations as discrete pose tokens in a learned latent space, and the Context-Guided Masked Transformer, which models token dependencies conditioned on both image features and 2D pose cues. This framework uses confidence-guided iterative sampling to refine reconstructions, producing realistic hand meshes under challenging conditions. MMHMR outperforms state-of-the-art methods, setting a new benchmark for 3D hand modeling in human-computer interaction, AR, and VR.
Appendix A Overview
  • Section B: Implementation Details
  • Section C: Ablation for VQ-MANO Pose Tokenizer
  • Section D: Generalization to Text-to-Mesh Generation
  • Section E: Effectiveness of Expectation-Approximated Differential Sampling
  • Section F: Confidence-Guided Masking
  • Section G: Pose Token Sampling Techniques
  • Section H: Impact of Control Signals for Hand Image Generation
  • Section J: Impact of Multi-Scale Features
  • Section K: Cross-Attention Layers in MMHMR
  • Section L: Qualitative Results
Appendix B Implementation Details
The implementation of MMHMR, developed using PyTorch, comprises two essential training phases: the VQ-MANO tokenizer and the context-guided masked transformer. These phases are meticulously designed to ensure accurate 3D hand mesh reconstruction while balancing computational efficiency and model robustness.
Context-Guided Masked Transformer. The second phase involves training the context-guided masked transformer, with the pose tokenizer frozen to leverage its pre-trained pose priors. This phase is dedicated to synthesizing pose tokens conditioned on input images and refining the 3D mesh reconstruction. Multi-resolution feature maps at 1×1$\times$1 × and 4×4$\times$4 × scales are used to capture both global and local contextual details, allowing the model to handle complex hand articulations and occlusions. The overall architecture of the system, including the Graph-Guided Pose Modeling (GGPM) and Context-Infused Masked Synthesizer, is depicted in Figure 5. The GGPM consists of two blocks ($xB$) of graph transformers to effectively model joint dependencies and ensure anatomical consistency. Meanwhile, the Context-Infused Masked Synthesizer employs four transformer layers ($xN$) to integrate multi-scale image features and refine pose token predictions through deformable cross-attention and token dependencies. The default number of iterations in Confidence-Guided Sampling is 5, which we use for the ablation study. The overall loss function integrates multiple objectives to guide the model toward robust reconstructions:
where $\mathcal{L}_{\text{mask}}$ minimizes errors in masked token predictions, $\mathcal{L}_{\text{MANO}}$ ensures consistency in MANO shape ($\beta$) and pose ($\theta$) parameters, $\mathcal{L}_{\text{3D}}$ aligns the predicted and ground-truth 3D joint positions, and $\mathcal{L}_{\text{2D}}$ preserves accurate 2D joint projections. The loss weights are configured as $\lambda_{\text{mask}}=1.0$, $\lambda_{\text{MANO}}=1.5\times 10^{-3}$ (with $\lambda_{\theta}=1\times 10^{-3}$ for pose and $\lambda_{\beta}=5\times 10^{-4}$ for shape), $\lambda_{\text{3D}}=5\times 10^{-2}$, and $\lambda_{\text{2D}}=1\times 10^{-2}$. This phase is trained using the Adam optimizer on NVIDIA RTX A6000 GPUs with a batch size of 48 and a learning rate of $1\times 10^{-5}$.
Mesh-Guided Control for Hand Image Generation. To extend the utility of the reconstructed 3D hand meshes, we integrate them as control signals into the publicly available ControlNet framework [66]. This integration enables the generation of high-quality hand images guided by the reconstructed meshes. The ControlNet models are trained on hand-cropped images resized to $256\times 256$, with a batch size of 64, using the Adam optimizer with a learning rate of $1\times 10^{-5}$. Training is conducted over 200 epochs on eight NVIDIA RTX A6000 GPUs. During training, image captions are generated using LLaVA-OneVision [36], a vision-language model capable of producing descriptive captions based on image content. Specifically, we query LLaVA-OneVision with prompts such as “What’s in the image and how does the hand pose look like?” to obtain concise descriptions focusing on the hand’s pose, including finger positions and hand orientation. For instance, in response to an image of a hand, LLaVA-OneVision might generate a caption like “A right hand with fingers slightly bent, palm facing upward.” These captions are then utilized as text guidance during the training of ControlNet. By combining these captions with the corresponding 3D hand meshes extracted using MMHMR, we enable the ControlNet framework to effectively learn a mapping between text, mesh signals, and high-quality hand image outputs. The generation pipeline involves two stages: training and inference. During training, MMHMR extracts 3D hand meshes from images, which are paired with text prompts to guide the ControlNet model. This process ensures that the generated images are both semantically aligned with the textual descriptions and anatomically consistent with the 3D hand meshes. At inference, the model requires only text prompts and mesh signals as inputs. The text prompts provide high-level semantic guidance, such as the desired pose or hand orientation, while the mesh signals ensure that the generated images maintain accurate anatomical and spatial features. This design enables the generation of realistic and contextually appropriate hand images, demonstrating the robustness and flexibility of the proposed approach. By leveraging both text-based and mesh-based guidance, the system achieves high fidelity in image generation, even in scenarios involving complex hand poses or occlusions.
B.1 Data Augmentation
In the initial training phase, the VQ-MANO module leverages prior knowledge of valid hand poses, serving as a critical foundation for the robust performance of the overall MMHMR pipeline. To deepen the model’s understanding of pose parameters, hand poses are systematically rotated across diverse angles, enabling it to effectively learn under varying orientations. In the subsequent training phase, the robustness of MMHMR is further enhanced through an extensive augmentation strategy applied to both input images and hand poses. These augmentations—such as scaling, rotations, random horizontal flips, and color jittering—introduce significant variability into the training data. By simulating real-world challenges like occlusions and incomplete pose information, these transformations prepare the model for complex, unpredictable scenarios. This comprehensive approach to data augmentation is a cornerstone of the training process, significantly improving the model’s ability to generalize and produce reliable, precise 3D hand mesh reconstructions across a wide range of conditions.
B.2 Camera Model
In the MMHMR pipeline, a simplified perspective camera model is employed to project 3D joints onto 2D coordinates, striking a balance between computational efficiency and accuracy. The camera parameters, collectively represented by $\Pi$, include a fixed focal length, an intrinsic matrix $K\in\mathbb{R}^{3\times 3}$, and a translation vector $T\in\mathbb{R}^{3}$. To streamline computations, the rotation matrix $R$ is replaced with the identity matrix $I_{3}$, further simplifying the model. The projection of 3D joints $J_{\text{3D}}$ onto 2D coordinates $J_{\text{2D}}$ is described as $J_{\text{2D}}=\Pi(J_{\text{3D}})$, where the operation encapsulates both the intrinsic parameters and the translation vector. This modeling approach reduces the parameter space, enabling computational efficiency while maintaining the accuracy required for robust 3D hand mesh reconstruction. By focusing on the most critical components, the model minimizes complexity without compromising performance.
Appendix C Ablation for VQ-MANO
Tables 8 and 9 summarize an ablation study on the Freihand [74] dataset, focusing on two key parameters: the number of pose tokens and the codebook size. Table 8 shows that increasing the number of pose tokens, while fixing the codebook size at $2048\times 256$, improves performance significantly, reducing PA-MPJPE from 1.01 mm to 0.41 mm and PA-MPVPE from 0.97 mm to 0.41 mm as tokens increase from 16 to 128. Table 9 highlights the effect of increasing the codebook size with a fixed token count of 64, showing a reduction in PA-MPJPE from 0.66 mm to 0.43 mm and PA-MPVPE from 0.65 mm to 0.44 mm as the size grows from $1024\times 256$ to $4096\times 256$. Notably, the codebook size has a stronger impact on performance than the number of pose tokens. The final configuration, with a codebook size of $2048\times 256$ and 64 tokens, balances efficiency and accuracy, achieving PA-MPJPE of 0.47 mm and PA-MPVPE of 0.44 mm. These results emphasize the importance of jointly optimizing these parameters for effective hand pose tokenization.
Number of Pose Tokens
Metric 16 32 64 128
PA-MPJPE 1.01 0.59 0.47 0.41
PA-MPVPE 0.97 0.57 0.44 0.41
Number of Codebook Size
Metric 1024 × 256 2048 × 128 2048 × 256 4096 × 256
PA-MPJPE 0.66 0.56 0.47 0.43
PA-MPVPE 0.65 0.58 0.44 0.44
Appendix D Generalization to Text-to-Mesh Generation
The MMHMR model is designed to be highly modular and adaptable, enabling functionality beyond image-conditioned tasks. To evaluate its generalization capability, we adapted MMHMR for text-to-mesh generation by replacing image-based conditioning with text guidance as shown in Figure 6. Leveraging the generative nature of MMHMR’s Masked Synthesizer, the model can generate diverse 3D meshes based on a single text input. This adaptability was further tested by removing the 2D pose guidance from MMHMR, making the model rely solely on text for mesh generation. The model was trained on the American Sign Language (ASL) dataset, following the same configuration as the image-conditioned MMHMR model but tailored for text-conditioned mesh synthesis. During testing, a top-5% probabilistic sampling strategy was applied, allowing the generation of multiple plausible meshes from a single text prompt. This experiment showcases the versatility and robustness of MMHMR in addressing novel tasks like text-to-mesh synthesis.
American Sign Language Dataset. The American Sign Language (ASL) dataset is a widely used resource for hand-based sign language recognition in English-speaking regions such as the United States and Canada. The dataset includes 26 letters of the alphabet, expressed using one-handed gestures, making it suitable for text-to-mesh experiments. For this study, we used dataset: the ASL alphabet dataset from Kaggle [1] for text to mesh generation.
Implementation Details and Results. Since the ASL datasets lack 2D or 3D annotations, such as poses, betas, or MANO parameters, we used our model MMHMR to generate pseudo-ground-truth (p-GT) annotations for the training images. These annotations were then employed to train MMHMR with text guidance, effectively transforming it into a text-to-mesh generator. Despite being derived annotations, the p-GT annotations were accurate enough to enable the generation of high-quality 3D meshes conditioned solely on text inputs. Moreover, we utilized CLIP [48] to extract text embeddings from the provided labels, enabling seamless integration of textual information into the generative pipeline. During testing, the model’s generative capability was demonstrated by producing multiple mesh outputs for a single text condition using a top-5% sampling strategy. This approach ensured diversity in the generated meshes while maintaining plausibility and fidelity to the text prompt. The generated meshes were quantitatively evaluated using standard metrics such as Hausdorff Distance and Chamfer Distance. Additionally, PA-MPVPE (measured in millimeters), which assesses vertex-to-vertex projection error, was introduced. These metrics were averaged across all generated meshes per text condition to ensure consistency. The results, presented in Table 10, indicate that MMHMR achieves high fidelity and consistency in text-to-mesh generation, with minimal deviations across samples. The quantitative evaluation affirms the precision of the generated meshes, while qualitative examples, illustrated in Figure 7, highlight the model’s ability to produce structurally accurate and visually realistic outputs for diverse text prompts. This experiment underscores MMHMR’s potential as a versatile generative framework for text-to-mesh applications.
Table 10: Quantitative Evaluation of Text-to-Mesh Generation using MMHMR’s Masked Synthesizer. Metrics include mean and standard deviation for Hausdorff Distance, Chamfer Distance, and PA-MPVPE (in mm). Results are averaged across all generated meshes per text prompt.
Metric Mean Standard Deviation
Hausdorff Distance 0.0221 0.0073
Chamfer Distance $9.73\times 10^{-5}$ $5.47\times 10^{-5}$
PA-MPVPE (mm) 12.2 3.1
Figure 6: Architecture for Text-to-Mesh Generation using MMHMR’s Context-Guided Masked Transformer. A text prompt $W$ is encoded using CLIP to generate text embeddings, which guide the mesh generation pipeline. The pipeline includes a Graph-Guided Pose Modeling module and a Context-Infused Masked Synthesizer, where expectation-approximated sampling selects pose tokens based on confidence scores. These pose tokens are processed through the VQ Encoder and decoded into mesh parameters $\theta^{\prime}$ using the VQ Decoder. This architecture highlights MMHMR’s ability to integrate textual context for accurate and generative 3D mesh synthesis.
Figure 7: Qualitative results of text-to-mesh generation for the American Sign Language (ASL) dataset. The generated meshes showcase the structural fidelity and diversity achieved by MMHMR’s Masked Synthesizer when conditioned solely on text inputs.
Appendix E Effectiveness of Expectation-Approximated Differential Sampling
@@
Appendix F Confidence-Guided Masking
@@
Appendix G Pose Token Sampling Techniques
The ablation study, illustrated in Table 12 and supported by Figure 9, highlights the crucial role of the Top-$k$-th token sampling strategy in MMHMR. Top-1st token sampling, which selects the most confident token prediction for each pose token, yields the most accurate results, with the lowest PA-MPJPE (7.0 mm for HO3Dv3 and 5.7 mm for FreiHAND) and PA-MPVPE (7.0 mm for HO3Dv3 and 5.5 mm for FreiHAND). This demonstrates the model’s ability to focus on high-confidence tokens, ensuring precise 3D hand mesh reconstructions. As $k$ increases, performance degrades significantly due to the inclusion of less confident tokens. For example, with Top-20th token sampling, PA-MPJPE increases to 10.57 mm on HO3Dv3 and 10.63 mm on FreiHAND, as noise and uncertainty introduced by broader sampling reduce the accuracy of the final pose estimation. This pattern is clearly reflected in Figure 9, where reconstructions using the 1st most confident token closely resemble the ground truth, while lower-ranked tokens (e.g., 100th or 1000th token) lead to unrealistic and distorted poses. These findings underscore the necessity of prioritizing high-confidence token predictions, as seen in the Top-1st token sampling strategy, to effectively address ambiguities in 2D-to-3D mapping. Broader sampling dilutes precision by introducing irrelevant or noisy tokens. The analysis reaffirms that MMHMR’s reconstruction fidelity hinges on the fine-grained selection of tokens, emphasizing the importance of a confidence-guided approach for robust and accurate performance.
Figure 9: Visualization of the impact of pose token confidence on MMHMR’s reconstruction accuracy. The figure compares 3D hand mesh reconstructions using the 1st, 7th, 10th, 100th, 250th, and 1000th most confident pose tokens. Reconstructions based on the 1st most confident token closely align with the ground truth, achieving high fidelity and structural accuracy. As confidence decreases (e.g., 100th to 1000th tokens), the reconstructions progressively deviate, with significant distortions and unrealistic poses evident. This highlights the critical role of selecting high-confidence tokens (Top-1) for accurate and robust 3D hand mesh reconstruction.
Table 12: Impact of Top-kth Sampling on MMHMR on HO3DV3 and FreiHAND dataset.
HO3Dv3 FreiHAND
Top-K-th PA-MPJPE PA-MPVPE PA-MPJPE ) PA-MPVPE
1 7.0 7.0 5.7 5.5
3 8.3 8.1 7.9 8.0
5 8.4 8.2 7.9 8.1
10 10.57 10.47 9.1 9.2
20 10.57 10.48 10.63 10.75
Appendix H Impact of Control Signals for Hand Image Generation
The role of control signals in guiding the generation of hand images using MMHMR’s Masked Synthesizer was evaluated through mesh-guided and pose-guided strategies. By leveraging meshes as control signals, the model demonstrated its ability to produce realistic and accurate hand images aligned with the underlying 3D structures. The results, presented in Table 13, show that mesh-guided control achieved better performance in terms of FID-H and KID-H, as well as higher hand detection confidence compared to 2D pose guidance. Notably, the combination of both mesh and 2D pose guidance further enhanced the model’s ability to generate high-quality images with fine structural details. The experiment highlighted that mesh guidance provides stronger structural constraints, leading to superior generation quality. The addition of 2D pose information complements the mesh signals by refining finer spatial details. These findings underline the versatility of MMHMR’s Masked Synthesizer in integrating multiple control modalities to achieve robust and visually coherent hand image generation, even in complex scenarios like hand-object interactions. Figures 10 and 11 demonstrate the effectiveness of mesh-guided and combined control strategies in generating realistic hand images. Mesh guidance provides strong structural constraints, while 2D pose signals refine spatial accuracy, ensuring robust and visually coherent outputs even in complex scenarios.
Table 13: Mesh-Guided Control for Hand Image Generation
HaMer Mesh Estimator FID-H $\downarrow$ KID-H $\downarrow$ Hand Det Conf. $\uparrow$
2D Pose Guidance 45.65 0.035 0.865
Mesh Guidance 40.23 0.027 0.912
Mesh + 2D Pose Guidance 40.01 0.026 0.910
Figure 10: Illustration of control strategies for hand image generation using Mesh-Guided ControlNet. The top figure part demonstrates the 2D Pose-Guided Control approach, where the 3D mesh estimator generates a 2D pose used to condition the ControlNet for generating the image. The bottom figure part depicts the combined 2D Pose + Mesh-Guided Control strategy, where both 2D pose and 3D mesh guidance are utilized for training, ensuring more accurate and realistic hand image generation. The figure highlights the distinction between inference and training processes, with prompts like ”The person is holding a can” guiding the generation pipeline.
Figure 11: Visualization of Mesh-Guided Control for Hand Image Generation. The left column illustrates 3D meshes used as control signals, while the right column displays the corresponding generated hand images. Mesh-guided generation accurately captures structural details and aligns with the intended hand poses and interactions, demonstrating the effectiveness of the control signals in MMHMR’s Masked Synthesizer.
Appendix I Impact of Pose Tokenizer on MMHMR
@@
The results presented in Table 14 demonstrate the critical influence of the Pose Tokenizer’s design on the performance of MMHMR. Increasing the codebook size from $1024\times 256$ to $2048\times 256$ yields significant improvements in both PA-MPJPE and MVE metrics across the HO3Dv3 and FreiHAND datasets. This indicates that a moderately larger codebook provides richer and more expressive pose representations, enabling better reconstruction of complex 3D hand poses. However, expanding the codebook further to $4096\times 256$ diminishes accuracy, suggesting that an overly large codebook introduces unnecessary complexity, making it harder for the model to generalize effectively.
Table 14: Impact of Codebook Size (Tokens = 96) on MMHMR.

### v2

@@ Appendix D In-the-Wild Reconstruction Evaluation
MaskHand Ours 61.0 87.1 94.8 62.1 90.2 95.0 59.3 88.3 94.4
Appendix E Occluded and Masked Hands Reconstruction
Figure 9 provides a qualitative, zero-shot comparison between MaskHand and the state-of-the-art HaMeR method on severely masked images (with hand regions masked around 90%). Despite significant occlusion, MaskHand consistently generates more plausible and anatomically accurate hand reconstructions than HaMeR. Specifically, MaskHand effectively synthesizes occluded regions, demonstrating robust generalization to previously unseen masked areas and preserving natural hand poses and orientations. In contrast, HaMeR exhibits noticeable reconstruction failures, inaccuracies, and unnatural poses, particularly under extreme masking conditions. These results highlight MaskHand’s superior capability in modeling uncertainty and synthesizing realistic meshes under severe occlusion, underscoring its robustness and practical effectiveness in challenging real-world scenarios.
Appendix F Impact of 2D Pose Context
We investigate how the accuracy of the 2D pose estimator influences MaskHand’s 3D reconstruction performance. In our main experiments, we utilize a lightweight OpenPose estimator OpenPose due to its computational efficiency and suitability for real-time applications. To quantify how improvements in 2D pose estimation may affect overall reconstruction quality, we conduct an additional analysis comparing OpenPose predictions against ground-truth 2D keypoints on the FreiHAND dataset. As shown in Table 10, using ground-truth 2D poses consistently yields better reconstruction metrics, notably improving both PA-MPJPE and PA-MPVPE. Although MaskHand achieves robust performance even with estimated keypoints, this analysis indicates that further advances in 2D pose estimation accuracy can directly enhance the quality of reconstructed 3D hand meshes.
Estimator PA-MPJPE PA-MPVPE F@5mm F@15mm
Ground Truth 5.2 5.1 0.834 0.993
2D OpenPose [7] Estimator 5.5 5.4 0.801 0.991
Table 10: Impact of 2D Pose Estimator on 3D Reconstruction Quality on Friehand dataset [77]
Appendix G Confidence-Aware Unconditional Mesh Generation
Methods APD(mm)↑ SI(%)↓
PCA 16.3 0.32
MaskHand (ours) 19.3 0.04
Table 11: Confidence-Aware Unconditional Mesh Generation
MaskHand enables confidence-aware unconditional 3D hand mesh generation by leveraging generative masked modeling and probabilistic sampling. We generate 2,000 hand meshes by setting the image condition to zero in the image encoder, ensuring that synthesis is driven entirely by the learned pose distribution. Using Top-100 sampling, MaskHand not only produces high-quality, diverse meshes but also quantifies the confidence of each generated hand configuration. This confidence estimation allows MaskHand to distinguish between physically plausible and invalid meshes, an advantage over diffusion-based HHMR [38], which lacks direct plausibility quantification.
In contrast, PCA-based generation, which samples from the MANO parameter space, produces structurally valid but limited and less diverse hand poses due to its restriction to the linear PCA subspace. The qualitative results (Figure 10) highlight MaskHand’s ability to synthesize highly articulated hand poses, while lower-confidence samples exhibit unnatural deformations, demonstrating its uncertainty quantification capability. Quantitatively, MaskHand achieves greater diversity than PCA, as reflected in a higher APD (19.3mm vs. 16.3mm), and generates more realistic meshes, reducing self-intersection (SI) from 0.32% to 0.04% (Table 11). These results confirm that MaskHand surpasses PCA in both diversity and realism, offering a principled approach to filtering implausible generations, making it a more reliable solution for unconditional 3D hand synthesis.
Figure 10: Confidence-Aware Unconditional Mesh Generation
Appendix H Effectiveness of Proposed MaskHand Components
The ablation study on HO3Dv3 (Table 12) highlights GAPR as the most critical component, ensuring joint dependencies and anatomical coherence. Removing the Upsampler and 2D Pose Context slightly reduces accuracy, affecting fine details and spatial cues. The full MaskHand model achieves the best results, demonstrating the importance of these components for high-precision 3D hand mesh recovery. Qualitative comparisons in Figure 11 further illustrate these effects, showing the impact of each component on reconstruction quality.
Method PA-MPJPE PA-MPVPE F@5mm AUCJ AUCV
@@ Appendix H Effectiveness of Proposed MaskHand Components
w/o. GAPR 7.3 7.3 0.645 0.853 0.854
MaskHand (Full) 7.0 7.0 0.663 0.860 0.860
Figure 11: Qualitative ablation study on component impact: Full model achieves highest accuracy, validating each component’s role.
Appendix I Masking Ratio during Training
The ablation study in Table 13 shows that a broader masking range $\gamma(\tau\in\mathcal{U}(0,0.7))$ achieves optimal results on HO3Dv3 and FreiHAND, with the lowest PA-MPVPE values of 7.0 and 5.5, respectively. This cosine-based masking strategy, where the model learns to reconstruct from partially masked sequences, enhances robustness in 3D hand reconstruction. Narrower masking ranges, such as $\gamma(\tau\in\mathcal{U}(0,0.3))$, increase error, highlighting the importance of challenging the model with broader masking for better generalization.
@@ Appendix I Masking Ratio during Training
Appendix J Effectiveness of Expectation-Approximated Differential Sampling
@@
Appendix K Confidence-Guided Masking
@@
Appendix L Impact of Pose Tokenizer on MaskHand
@@
The results presented in Table 15 demonstrate the critical influence of the Pose Tokenizer’s design on the performance of MaskHand. Increasing the codebook size from $1024\times 256$ to $2048\times 256$ yields significant improvements in both PA-MPJPE and MVE metrics across the HO3Dv3 and FreiHAND datasets. This indicates that a moderately larger codebook provides richer and more expressive pose representations, enabling better reconstruction of complex 3D hand poses. However, expanding the codebook further to $4096\times 256$ diminishes accuracy, suggesting that an overly large codebook introduces unnecessary complexity, making it harder for the model to generalize effectively.
Table 15: Impact of Codebook Size (Tokens = 96) on MaskHand.